OpenStudy (anonymous):

1.2x + 3.8 = 0.8x + 1.6

OpenStudy (anonymous):

x=-11/2

OpenStudy (anonymous):

your answer is wrong

OpenStudy (anonymous):

1.2x+3.8=0.8x+1.6 Since 0.8x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 0.8x from both sides. 1.2x+3.8-0.8x=1.6 Since 1.2x and -0.8x are like terms, add -0.8x to 1.2x to get 0.4x. 0.4x+3.8=1.6 Since 3.8 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3.8 from both sides. 0.4x=-3.8+1.6 Add 1.6 to -3.8 to get -2.2. 0.4x=-2.2 Divide each term in the equation by 0.4. (0.4x)/(0.4)=-(2.2)/(0.4) Cancel the common factor of 0.4. x=-(2.2)/(0.4) Divide -2.2 by 0.4 to get -5.5. x=-5.5

OpenStudy (anonymous):

Just because you can't divide -11/2 doesn't mean that it's wrong, it means you need to look before leaping to conclusions.